# Rank jumps and Multisections of elliptic fibrations on K3 surfaces

### v1

@@ 5.1. Rank jump properties for codimension one subfamilies
First we observe that if a K3 surface has an elliptic fibration and Picard number 4, its Picard group is $U\oplus\Gamma$ with basis $F,\mathcal{O},b_{3},b_{4}$, and with respect to the fibration whose fiber is $F$ one of the following three cases appears:
@@
    $\Gamma$ does not contain vectors of self-intersection $-2$: in this case the elliptic fibration has no reducible fibers and the Mordell–Weil group has rank 2;
@@
    $\Gamma$ contains a class of self-intersection $-2$ and the root lattice of $\Gamma$ has rank 1: in this case there is exactly one reducible fiber, which has two components, and the Mordell–Weil group has rank 1;
@@
In case $b)$ there are two possibilities, namely that the generator of the Mordell–Weil group intersects the reducible fiber as the section $\mathcal{O}$, or that it intersects in the other component. If the former holds, the Mordell–Weil lattice (see [ScS10, Section 11] for definition) has an integer discriminant and the lattice $\Gamma$ is isometric to $\langle-2\rangle\oplus\langle-2\gamma\rangle$ with $\gamma>1$; in this case the Mordell–Weil group is $\mathbb{Z}$ and the classes of the sections are $P_{n}=\gamma n^{2}F+\mathcal{O}+nb_{4}$. If the latter holds, the Mordell–Weil lattice has a non-integer discriminant and the lattice $\Gamma$ is isometric to $\left[\begin{array}[]{cc}-2&-1\\ -1&-2\gamma\end{array}\right]$ with $\gamma>1$; in this case the Mordell–Weil group is $\mathbb{Z}$; the classes of the sections corresponding to even integers are $P_{-2n}=(4\gamma-1)n^{2}F+\mathcal{O}+nb_{3}-2nb_{4}$ and the ones corresponding to odd integers are $P_{-2n-1}=\left((4\gamma-1)n^{2}+(4\gamma-1)n+\gamma\right)F+\mathcal{O}+nb_{3}+(-2n-1)b_{4}$.
@@
This follows as in Theorem 5.3, by applying Remark 5.5 to the class $B=\frac{d+3}{2}F+2\mathcal{O}+b_{3}$, which is such that $B^{2}=-2$, $BF=2>0$ and $B\mathcal{O}\equiv 1\mod 2$. It is effective and we now show that it cannot be written as the sum with non-negative coefficients of two sections and components of fibers. The result is true both if $c=1$ and hence the Mordell–Weil group of the fibration has rank 1 and if $c>1$ and hence the Mordell–Weil group of the fibration has rank 2. We give the details only in the case $c>1$.
If there exist $r,x,y\in\mathbb{N}$ such that $B=rF+xP_{m,n}+yP_{p,q}$, then $x+y=2$, $xm+yp=1$, $xn+yq=0$. This implies that $x=y=1$, $p=1-m$ and $q=-n$. By (5.1), $P_{m,n}+P_{1-m,-n}=(2dm^{2}+2cn^{2}-2md+d)F+2\mathcal{O}+b_{3}$ with $d\geq 5$. The coefficient of $F$ is larger than $(d+3)/2$ for every choice of $m$ and $n$, hence $r$ is forced to be a negative number, which is not acceptable. ∎
@@
$\ U\oplus\left[\begin{array}[]{rr}-6&-1\\ -1&-6\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-4&0\\ 0&-10\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-4&0\\ 0&-6\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-2&-1\\ -1&-10\end{array}\right]\\ &&&\\ U\oplus\left[\begin{array}[]{rr}-2&-1\\ -1&-6\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-2&-1\\ -1&-4\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-2&1\\ 1&-2\end{array}\right]&U\oplus\left[\begin{array}[]{rr}-2&0\\ 0&-2\end{array}\right]\\ &&&\\ \hline\cr\end{array}$
Table 1. Intersection form for Picard lattices of K3s with a unique elliptic fibration known to have the potentially rank jump property
@@
We first identify an effective class $B$ such that $BF=2$ and $B^{2}=-2$, and then we prove that it is supported on an irreducible curve (in particular it is not the sum of two sections and possibly some components of fibers). This guarantee that it really corresponds to a bisection and satisfies Remark 5.5.
For all the lattices considered here, one can prove that there exists a basis $\{b_{3},b_{4}\}$ of $\Gamma$ such that the bilinear form is $\left[\begin{array}[]{rr}-2\alpha&\beta\\ \beta&-2\gamma\end{array}\right]$, with $\gamma>0$, and $\alpha\geq 5$, $\alpha\equiv 1\mod 2$. In particular $\left[\begin{array}[]{cc}-6&-1\\ -1&-6\end{array}\right]\simeq\left[\begin{array}[]{cc}-10&-5\\ -5&-6\end{array}\right]$ and $\left[\begin{array}[]{rr}-4&0\\ 0&-6\end{array}\right]\simeq\left[\begin{array}[]{cc}-10&-4\\ -4&-4\end{array}\right]$.
@@
which implies $y+z=2$, $ym+zp=1$, $yn+zq=0$, by comparing the coefficients of $\mathcal{O}$, $b_{3}$ and $b_{4}$ respectively. In particular, one obtains that $y=z=1$, $p=1-m$, $q=-n$ and one concludes since $x=\frac{\alpha+3}{2}-\left(2\alpha m^{2}-2\beta mn+2\gamma n^{2}+\alpha-2m\alpha+\beta n\right)$ is a negative number for any possible choice of $(m,n)\in\mathbb{Z}\times\mathbb{Z}$. Therefore $B$ is irreducible.
Second case: the Mordell Weil group has rank 1 (case $b)$). The cases which appear in Table 1 are $\Gamma=\left[\begin{array}[]{cc}-2&-1\\ -1&-2\gamma\end{array}\right]$ with $\gamma=2,3,5$. For $\gamma=2,3,5$, consider the class $B=16F+\mathcal{O}+b_{3}-4b_{4}$, $B=24F+\mathcal{O}+b_{3}-4b_{4}$ and $B=4F+\mathcal{O}-b_{4}$, respectively. It has the required properties (i.e. $BF=2$, $B^{2}=-2$) and cannot be written as $rF+xb_{3}+yP_{a}+zP_{b}$ with $y+z=2$ and $r$ and $x$ non negative.
@@
If $\Gamma\simeq\left[\begin{array}[]{rr}-4&-2\\ -2&-4\end{array}\right]$, then it is not possible to find in $\Gamma$ a class with self-intersection $-2d$ such that $d$ is odd, hence it is not possible to show the existence of a smooth rational bisection with the same method applied before. Moreover, since [GM24] guarantees that there are no more than one genus one fibration on the surface, the elliptic fibration does not admit a smooth genus 1 bisection.
If $\Gamma\simeq\left[\begin{array}[]{rr}-2&0\\ 0&-4\end{array}\right]$ in order to find a rational bisection, we need a class $B=xF+y\mathcal{O}+zb_{3}+wb_{4}$ such that $B^{2}=-2$, $BF=2$, $0\leq Bb_{3}\leq 2$ (since $b_{3}$ is identified with the class of an irreducible component of a reducible fiber). This forces $y=2$, $z=-1$ and hence $x=2+w^{2}$. So $B=(2+w^{2})F+2\mathcal{O}+b_{3}+wb_{4}$, which is equal to $2F+b_{3}+2P_{w/2}$ if $w$ is even and to $b_{3}+P_{v}+P_{v-1}$ if $w=2v-1$. In both the cases $B$ corresponds to the union of sections and components of fibers, and so it does not correspond to a rational bisection. So for this choice of $\Gamma$ the fibration does not admit a smooth bisection of genus either 0 or 1.
@@
6. Example
In this section, we describe several examples for which we prove the potential rank jump property and/or the rank jump property on very explicit K3 surfaces with Picard number 3, for which we give an equation. We give two examples of elliptic K3 surfaces with Picard rank 3 over $k=\mathbb{Q}(\sqrt{5})$, realized as a double cover of a nodal sextic with principal tangents defined only over a degree 2 extension. By Example 6.1, the bisection above the node is $\mathbb{P}^{1}_{\mathbb{Q}}$. By Example 6.2, it is a conic without $k$-points. Therefore, by the proof of Theorem 5.3, the former admits the rank jump property over $k$, while the latter admits the property after a suitable degree 2 extension of $k$.
@@ 6. Example
Then $C$ has a unique singularity, namely $p=(1:0:0)$, which is of type $A_{1}$. It admits a tritangent line, namely $L:x=0$.
Let $X$ be the K3 double cover of $\mathbb{P}^{2}$ branched over $C$, then the geometric Picard number of $X$ is at least 3. Indeed, its Picard lattice contains $B$, the rational curve above the exceptional divisor given by the desingularization of $p$, a genus 1 class $F$, given by the pull-back of any line through $p$, and the class $\mathcal{O}$ given by one of the two irreducible rational curves above the tritangent line $L$. Since $p\not\in L$, their intersection matrix is a follows
@@
and in particular $x=B\mathcal{O}=0$, which implies $d=5$, i.e., the Picard group contains a lattice isometric to $\Lambda_{5}=U\oplus\langle-10\rangle$.
We now prove that $\rho(X)=3$, which will imply that $\operatorname{Pic}(X)\simeq\Lambda_{5}$ and that $L$ is the unique tritangent to the curve $C$.
@@
In our case, a computation of the characteristic polynomial of the Frobenius for ordinary primes such as $7,11,23,31$ and 37, done with the help of Magma ([Ma]), yields a factor of degree 18 that is not cyclotomic. Hence $\rho\leq 4$.
@@
For the prime $7$, the sextic $C$ has two singularities of type $A_{1}$, namely $p=(1:0:0)$ and $q=(1:5:1)$. The exceptional divisor above $q$ is a smooth rational curve. So the Picard group for $p=7$ is the lattice $\Lambda_{5}\oplus\langle-2\rangle$, where the last summand is spanned by the exceptional divisor over $q$. If this curve lifted over $\mathbb{Q}$, then the Picard group of $X$ over $\mathbb{Q}$ would be $\Lambda_{5}\oplus\langle-2\rangle$ and in particular the elliptic fibration induced by the pencil of lines through $p$ would have a fiber of type $I_{2}$. But one can directly check on the equation of the fibration, that it does not admit any reducible fiber over $\mathbb{Q}$. Therefore the class which is defined for $p=7$ as exceptional divisor over $q$ does not lift over $\mathbb{Q}$ and this implies that $\rho=3$ over $\mathbb{Q}$.
Finally, we verify that the smooth bisection $B$ is saliently ramified. Recall that the unique elliptic fibration on $X$ is given by the pull-back of lines through $p$. Moreover, the restriction map $\pi_{B}:B\rightarrow\mathbb{P}^{1}$ is branched over the fibers that are pull-backs of the two lines that are the principal tangents of $C$ at the node $p$, namely $y=2iz$ and $y=-2iz$. Since these intersect $C$ transversally outside $p$, they pull-back to smooth fibers of $\pi$.
@@
The elliptic fibration has no reducible fibers.
@@
We observe that since the elliptic fibration has infinitely many sections, there are infinitely many irreducible curves in $\mathbb{P}^{2}$ that split in the double cover. The section $S_{n}$ comes from the splitting of a curve of degree $110n^{2}-10n+$ 1 with a singular point of multiplicity $10n^{2}+10n$ in $p$.
@@
Then, as before, $C$ has a unique singularity, namely $p=(1:0:0)$, which is of type $A_{1}$. It admits a tritangent line, namely $L:x=0$, which turns out to be unique, once one proved as above that $\rho=3$.
@@
Then, by [NvL24, Thm. 3.3], $X$ is a K3 surface with Picard number 3. Its Picard lattice is generated by the classes of $H,L_{0},L_{1}$, where $H$ corresponds to a hyperplane section, $L_{0}$ is the class of the line given by $x=z=0$ and $L_{1}$, the class of $x-y=z-w=0$. Moreover, it is $U\oplus\langle-20\rangle$-polarized. In fact, the divisor $D=6H-7L_{0}+3L_{1}$ is perpendicular to a copy of $U$ and has self-intersection $-20$.
@@
Consider the line $L_{0}$ and the pencil of planes $x=tz$ through that line. The intersection of the quartic with $x=tz$ contains the line $L_{0}$; the residual cubic is the fiber of the elliptic fibration $\pi_{0}$ over the point $t$. One can directly check that for $t=0$ the intersection between the line and the residual cubic consist of the points $(0:0:0:1)$ and $(0:1:0:0)$ and the first one has multiplicity two. In particular the trisection (which corresponds to $L_{0}$) branches on the fiber over $t=0$. Nevertheless, one can also check that the intersection $X\cap\{x=0\}$ consists of the line $L_{0}$ and of a smooth cubic, so the trisection $L_{0}$ branches over a smooth fiber of the fibration, i.e. it is saliently ramified.

### v2

@@ 5.1. Rank jump properties for codimension one subfamilies
First we observe that if a K3 surface has an elliptic fibration and Picard number 4, its Picard group is $U\oplus\Gamma$ with base $F,\mathcal{O},b_{3},b_{4}$, and with respect to the fibration whose fiber is $F$ one of the following three cases appears:
@@
    $\Gamma$ does not contain vectors of self-intersection 2-2: in this case the elliptic fibration has no reducible fibers and the Mordell–Weil group has rank 2;
@@
    $\Gamma$ contains a class of self-intersection 2-2 and the root lattice of $\Gamma$ has rank 1: in this case there is exactly one reducible fiber, which has two components, and the Mordell–Weil group has rank 1;
@@
In case $b)$ there are two possibilities, namely that the generator of the Mordell–Weil group intersects the same component of the reducible fiber as the section $\mathcal{O}$, or that it intersects in the other component. If the former holds, the Mordell–Weil lattice (see [ScS10, Section 11] for definition) has an integer discriminant and the lattice $\Gamma$ is isometric to $\langle-2\rangle\oplus\langle-2\gamma\rangle$ with $\gamma>1$; in this case the Mordell–Weil group is $\mathbb{Z}$ and the classes of the sections are $P_{n}=\gamma n^{2}F+\mathcal{O}+nb_{4}$. If the latter holds, the Mordell–Weil lattice has a non-integer discriminant and the lattice $\Gamma$ is isometric to $\left[\begin{array}[]{cc}-2&-1\\ -1&-2\gamma\end{array}\right]$ with $\gamma>1$; in this case the Mordell–Weil group is $\mathbb{Z}$; the classes of the sections corresponding to even integers are $P_{-2n}=(4\gamma-1)n^{2}F+\mathcal{O}+nb_{3}-2nb_{4}$ and the ones corresponding to odd integers are $P_{-2n-1}=\left((4\gamma-1)n^{2}+(4\gamma-1)n+\gamma\right)F+\mathcal{O}+nb_{3}+(-2n-1)b_{4}$.
@@
This follows as in Theorem 5.3, by applying Remark 5.5 to the class $B=\frac{d+3}{2}F+2\mathcal{O}+b_{3}$, which is such that $B^{2}=-2$, $BF=2>0$ and $B\mathcal{O}\equiv 1\mod 2$. It is effective and, in the light of Remark 5.5, we now show that it cannot be written as the sum with non-negative coefficients of sections and components of fibers. In what follows, we give the details only in the case $c>1$, which corresponds to case a) in 5.1 above. In that case, the Mordell–Weil rank is 2 and all fibers are irreducible. Assume, by contradiction, that $B$ can be written as a sum of fibers and sections. This implies that there are $r,x,y\in\mathbb{N}$ such that $B=rF+xP_{m,n}+yP_{p,q}$. By the condition $BF=2$, it follows $x+y=2$; by the comparison of the coefficients of $b_{3}$ (resp. $b_{4}$) in $B=\frac{d+3}{2}F+2\mathcal{O}+b_{3}$ and in $B=rF+xP_{m,n}+yP_{p,q}$, one obtains $xm+yp=1$ (resp. $xn+yq=0$). Therefore $x=y=1$, p=1mp=1-m and q=nq=-n. By (5.1), $P_{m,n}+P_{1-m,-n}=(2dm^{2}+2cn^{2}-2md+d)F+2\mathcal{O}+b_{3}$ with $d\geq 5$. The coefficient of $F$ is larger than $(d+3)/2$ for every choice of $m$ and $n$, hence $r$ is forced to be a negative number, yielding the desired contradiction.
The case where $c=1$, implies the existence of a vector of self-intersection 2-2 which is orthogonal to the class of a fiber, yielding a fiber component. This corresponds to case b) in 5.1. One obtains a contradiction in a similar way: one considers the class $B=\frac{d+3}{2}F+2\mathcal{O}+b_{3}$ and one shows that it is not possible that $B$ can be also expressed as $rF+xP_{m}+yP_{n}+zb_{3}$ with $r,x,y,z\geq 0$. One has to be careful since the expressions of $P_{m}$ and $P_{n}$ in terms of the basis $\{F,\mathcal{O},b_{3},b_{4}\}$, depends on the parity of $m$ and $n$, so one has to consider all the possible cases (i.e. $m$ and $n$ are both even, are both odd, one is odd and one is even). ∎
@@
[…]
Table 1. Intersection form for Picard lattices of rank 4 of K3s with a unique elliptic fibration known to have the potentially rank jump property
@@
We first identify an effective class $B$ such that $BF=2$ and $B^{2}=-2$, and then we prove that it is supported on an irreducible curve (in particular it is not the sum of two sections and possibly some fibers). This guarantees that it indeed corresponds to an irreducible bisection and satisfies Remark 5.5.
For all the lattices considered here, one can prove that there exists a base $\{b_{3},b_{4}\}$ of $\Gamma$ such that the bilinear form is $\left[\begin{array}[]{rr}-2\alpha&\beta\\ \beta&-2\gamma\end{array}\right]$, with $\gamma>0$, and $\alpha\geq 5$, $\alpha\equiv 1\mod 2$. In particular $\left[\begin{array}[]{cc}-6&-1\\ -1&-6\end{array}\right]\simeq\left[\begin{array}[]{cc}-10&-5\\ -5&-6\end{array}\right]$ and $\left[\begin{array}[]{rr}-4&0\\ 0&-6\end{array}\right]\simeq\left[\begin{array}[]{cc}-10&-4\\ -4&-4\end{array}\right]$.
@@
which implies $y+z=2$, $ym+zp=1$, $yn+zq=0$, by comparing the coefficients of $\mathcal{O}$, $b_{3}$ and $b_{4}$ respectively. In particular, one obtains that $y=z=1$, p=1mp=1-m, q=nq=-n and one concludes since $x=\frac{\alpha+3}{2}-\left(2\alpha m^{2}-2\beta mn+2\gamma n^{2}+\alpha-2m\alpha+\beta n\right)$ is a negative number for any possible choice of $(m,n)\in\mathbb{Z}\times\mathbb{Z}$. Therefore $B$ is irreducible.
Second case: the Mordell Weil group has rank 1 (case $b)$). The cases which appear in Table 1 are $\Gamma=\left[\begin{array}[]{cc}-2&-1\\ -1&-2\gamma\end{array}\right]$ with $\gamma=2,3,5$. For $\gamma=2,3,5$, consider the class $B=16F+2\mathcal{O}+b_{3}-4b_{4}$, $B=24F+2\mathcal{O}+b_{3}-4b_{4}$ and $B=4F+2\mathcal{O}-b_{4}$, respectively. It has the required properties (i.e. $BF=2$, $B^{2}=-2$) and cannot be written as $rF+xb_{3}+yP_{a}+zP_{b}$ with $y+z=2$ and $r$ and $x$ non negative, where $b_{3}$ is the class of an irreducible component of the reducible fiber and $P_{a}$ and $P_{b}$ are sections.
@@
If $\Gamma\simeq\left[\begin{array}[]{rr}-4&-2\\ -2&-4\end{array}\right]$, then it is not possible to find in $\Gamma$ a class with self-intersection 2d-2d such that $d$ is odd, hence it is not possible to show the existence of a smooth rational bisection with the same method applied before. Moreover, since [GM24] guarantees that there are no more than one genus 1 fibration on the surface, the elliptic fibration does not admit a smooth genus 1 bisection.
If $\Gamma\simeq\left[\begin{array}[]{rr}-2&0\\ 0&-4\end{array}\right]$ in order to find a rational bisection, we need a class $B=xF+y\mathcal{O}+zb_{3}+wb_{4}$ such that $B^{2}=-2$, $BF=2$, $0\leq Bb_{3}\leq 2$ (since $b_{3}$ is identified with the class of an irreducible component of a reducible fiber). This forces $y=2$, z=1z=-1 and hence $x=2+w^{2}$. So $B=(2+w^{2})F+2\mathcal{O}+b_{3}+wb_{4}$, which is equal to $2F+b_{3}+2P_{w/2}$ if $w$ is even and to $b_{3}+P_{v}+P_{v-1}$ if w=2v1w=2v-1. In both the cases $B$ corresponds to the union of sections and components of fibers, and so it does not correspond to a smooth rational bisection. So for this choice of $\Gamma$ the fibration does not admit a smooth bisection of genus either 0 or 1.
@@
6. Examples
In this section, we describe several examples for which we prove the potential rank jump property and/or the rank jump property on very explicit K3 surfaces with Picard number 3, for which we give an equation. We give two examples of elliptic K3 surfaces with Picard rank 3 over $k=\mathbb{Q}(\sqrt{5})$, realized as a double cover of a nodal sextic with principal tangents defined only over a degree 2 extension. In Example 6.1, the bisection above the node is $\mathbb{P}^{1}_{\mathbb{Q}}$. In Example 6.2, it is a conic without $k$-points. Therefore, by the proof of Theorem 5.3, the former admits the rank jump property over $k$, while the latter admits the property after a suitable degree 2 extension of $k$.
@@ 6. Examples
Then $C$ has a unique singularity, namely $p=(1:0:0)$, which is of type $A_{1}$. It admits a bitangent line $L:x=0$. As this line intersects the curve in only 2 points, namely $(0:1:0)$ and $(0:0:1)$, with intersection multiplicities 2 and 4, respectively, it can be considered as a degenerate tritangent line. In particular, it splits in the double cover $g:X\rightarrow\mathbb{P}^{2}$ branched on $C$. Indeed, if a curve $D$ on a surface $Y$ intersects the branch locus of a double cover $g:X\rightarrow Y$ with an with even multiplicity in a point $q$, then its inverse image $g^{-1}(D)\subset X$ is singular in $g^{-1}(q)$ but does not ramify in $g^{-1}(q)$ (because of the splitting of the local equation of the double cover). So, for the double cover $g:X\rightarrow\mathbb{P}^{2}$ branched on $C$, the inverse image of the line $L$ does not contain ramification points but has two singular points (which are $g^{-1}(0:1:0)$ and $g^{-1}(0:0:1)$). In particular, since $L$ is rational and so simply connected, the inverse image $g^{-1}(L)$ consists of two rational curves meeting in two points (one with multiplicity 2).
Let $X$ be the K3 double cover of $\mathbb{P}^{2}$ branched over $C$, then the geometric Picard number of $X$ is at least 3. Indeed, its Picard lattice contains $B$, the rational curve above the exceptional divisor given by the desingularization of $p$, a genus 1 class $F$, given by the pull-back of any line through $p$, and the class $\mathcal{O}$ given by one of the two irreducible rational curves above the line $L$. Since $p\not\in L$, their intersection matrix is as follows
@@
and, with the same notation of Section 3.3, we have $x=B\mathcal{O}=0$, which implies $d=5$, i.e., the Picard group contains a lattice isometric to $\Lambda_{5}=U\oplus\langle-10\rangle$.
We now prove that $\rho(X)=3$, which will imply that $\operatorname{Pic}(X)\simeq\Lambda_{5}$ and that $L$ is the unique line in $\mathbb{P}^{2}$ which splits in the double cover.
@@
In our case, a computation of the characteristic polynomial of the Frobenius for ordinary primes such as $7,11,23,31$ and 37, done with the help of Magma ([Ma]), yields a factor of degree 18 that is not cyclotomic. Since the characteristic polynomial of Frobenius has degree $b_{2}=22$, we obtain $\rho\leq 22-18=4$.
@@
For the prime $7$, the sextic $C$ has two singularities of type $A_{1}$, namely $p=(1:0:0)$ and $q=(1:5:1)$. The exceptional divisor above $q$ is a smooth rational curve. So the Picard group for $p=7$ is the lattice $\Lambda_{5}\oplus\langle-2\rangle$, where the last summand is spanned by the exceptional divisor over $q$. If this curve lifted over $\mathbb{Q}$, then the Picard group of $X$ over $\mathbb{Q}$ would be $\Lambda_{5}\oplus\langle-2\rangle$ and in particular the elliptic fibration induced by the pencil of lines through $p$ would have a fiber of type $I_{2}$. But one can directly check on the equation of the fibration (see, 6.1), that it does not admit any reducible fiber over $\mathbb{Q}$. Therefore the class which is defined for $p=7$ as exceptional divisor over $q$ does not lift over $\mathbb{Q}$ and this implies that $\rho=3$ over $\mathbb{Q}$.
Finally, we verify that the smooth bisection $B$ is saliently ramified. Recall that the unique elliptic fibration on $X$ is given by the pull-back of lines through $p$. Moreover, the restriction map $\pi_{B}:B\rightarrow\mathbb{P}^{1}$ is branched over the fibers that are pull-backs of the two lines that are the principal tangents of $C$ at the node $p$, namely $y=2iz$ and y=2izy=-2iz. Since these intersect $C$ transversally outside $p$, they pull-back to smooth fibers of $\pi$.
@@
The elliptic fibration has no reducible fibers, as can be checked by the computation of its discriminant.
@@
We observe that since the elliptic fibration has infinitely many sections, there are infinitely many irreducible curves in $\mathbb{P}^{2}$ that split in the double cover. The section $S_{n}$ comes from the splitting of a curve of degree $10n^{2}-10n+1$ with a singular point of multiplicity $10n^{2}+10n$ in $p$.
@@
Then, as before, $C$ has a unique singularity, namely $p=(1:0:0)$, which is of type $A_{1}$. It admits a degenerate tritangent line, namely $L:x=0$, which turns out to be unique, once one proved as above that $\rho=3$.
@@
Then, by [NvL24, Thm. 3.3], $X$ is a K3 surface with Picard number 3. Its Picard lattice is generated by the classes of $H,L_{0},L_{1}$, where $H$ corresponds to a hyperplane section, $L_{0}$ is the class of the line given by $x=z=0$ and $L_{1}$, the class of xy=zw=0x-y=z-w=0. Moreover, it is $U\oplus\langle-20\rangle$-polarized. In fact, the divisor $D=6H-7L_{0}+3L_{1}$ is perpendicular to a copy of $U$ and has self-intersection 20-20.
@@
Consider the line $L_{0}$ and the pencil of planes $x=tz$ through that line. The intersection of the quartic with $x=tz$ contains the line $L_{0}$; the residual cubic is the fiber of the elliptic fibration $\pi_{0}$ over the point $t$. One can directly check that for $t=0$ the intersection between the line and the residual cubic consists of the points $(0:0:0:1)$ and $(0:1:0:0)$ and the first one has multiplicity two. In particular, the trisection (which corresponds to $L_{0}$) branches on the fiber over $t=0$. Nevertheless, one can also check that the intersection $X\cap\{x=0\}$ consists of the line $L_{0}$ and of a smooth cubic, so the trisection $L_{0}$ branches over a smooth fiber of the fibration, i.e. it is saliently ramified.